# An Explicit M1 Radiation-hydrodynamics Scheme for Three-dimensional Protostellar Evolution

### v1

@@ 4.1 Hybrid radiative transfer: RSLA and Non-RSLA components
As discussed in Section 3.1, applying the RSLA violates physical energy conservation, making it difficult to accurately follow the time evolution of optically thick systems, especially when the radiation energy dominates over the gas internal energy. However, in very optically thick regions, radiation propagates slowly via diffusion, and therefore it is possible to evolve the radiation without applying the RSLA while maintaining a reasonably large time step. Indeed, considering the diffusion approximation given in Equation (13), von Neumann stability analysis requires the time step to satisfy
@@
Motivated by this, we introduce an additional radiation energy component, denoted as $E^{\mathrm{N}}_{\mathrm{rad}}$, which is evolved without the RSLA and dominates in regions where $\chi_{\mathrm{0R}}\geq\chi_{\mathrm{crit}}$. Hereafter, we refer to $E^{\mathrm{N}}_{\mathrm{rad}}$ as the non-RSLA component and to $E^{\mathrm{R}}_{\mathrm{rad}}$ as the RSLA component. As noted in Section 2.3, the conserved quantity of the RSLA component in the gas-radiation interaction is $E^{\mathrm{R}}_{\mathrm{rad}}$, which corresponds to the radiation energy density divided by the reduction factor $f_{c}$ as shown in Equation (35). In contrast, the radiation energy of the non-RSLA component $E^{\mathrm{N}}_{\mathrm{rad}}$ is itself a conserved quantity, as it evolves with the full speed of light. Therefore, in the region where $\chi_{\mathrm{0R}}\geq\chi_{\mathrm{crit}}$, we can accurately follow the evolution of optically thick systems while conserving energy.
@@
during the calculation. The timing of this redistribution can be chosen appropriately depending on the specific implementation of the code (see Section 5).
@@
where $\bm{G}^{\mathrm{R}}_{\mathrm{S}}+\bm{G}^{\mathrm{R}}_{\mathrm{T}}$ and $\bm{G}^{\mathrm{N}}$ are spatial components of the radiation four-force density of RSLA and non-RSLA components, respectively. From Equations (53) and (55), we can see that the $E_{\mathrm{gas}}+E^{\mathrm{con}}_{\mathrm{rad}}$ is conserved during the gas-radiation interaction. Note that we can solve Equations (53)–(59) in combination with Equations (1), (16), (43), (46), and (49).
It should be noted that energy conservation may locally break down in optically thick systems if they contain regions where $\tau_{\mathrm{cell}}<2/(3f_{c})$, since the non-RSLA component $E^{\mathrm{R}}_{\mathrm{rad}}$ dominates in such regions. This issue can become significant in extremely high-resolution simulations, where $\tau_{\mathrm{cell}}$ may fall below $2/(3f_{c})$ even within stellar interiors. However, in typical RHD star formation simulations, where the Jeans length is resolved by several tens of cells, $\tau_{\mathrm{cell}}$ remains sufficiently large throughout most of the stellar interior, and regions with $\tau_{\mathrm{cell}}<2/(3f_{c})$ are confined to narrow layers near the stellar surface that cannot be adequately resolved. Thus, due to the intrinsic limitation of resolution, our scheme cannot capture the detailed structure of these layers. Nevertheless, our scheme is designed to correctly calculate the net radiation energy transfer across the stellar surface, as described in Section 4.2.
4.2 Radiation distribution based on neighboring cell information
We distribute the conserved quantity $E^{\mathrm{con}}_{\mathrm{rad}}$ among $E^{\mathrm{R}}_{\mathrm{rad,S}}$, $E^{\mathrm{R}}_{\mathrm{rad,T}}$, and $E^{\mathrm{N}}_{\mathrm{rad}}$ based on the optical depth per cell $\tau_{\mathrm{cell}}$, as described in Equations (16) and (49). However, as mentioned in Section 3.2, this approach fails to capture the leakage of radiation from optically thick cells into neighboring optically thin cells, because the radiation energy in an optically thick cell is stored in either $E^{\mathrm{R}}_{\mathrm{rad,T}}$ or $E^{\mathrm{N}}_{\mathrm{rad}}$, neither of which is allowed to freely leave the cell. As a result, we cannot capture the outgoing radiation flux from optically thick cell surfaces. This issue is particularly relevant to the narrow layers near the stellar surface, where the gas density and opacity rise steeply.
@@ 4.2 Radiation distribution based on neighboring cell information
which approximately reproduces the physically expected surface flux. We describe in Section 5 the details of when and how the optical depths of neighboring cells are incorporated into the calculation, together with the code implementation.
4.3 Spectral reconstruction and opacity estimation from mean photon energy
As mentioned in Section 3.3, to accurately capture the interaction between the gas and radiation, we need to know the radiation spectrum. However, the radiation energy density alone does not uniquely determine the spectral shape. To address this issue, we evolve not only the radiation energy densities, $E^{\mathrm{R}}_{\mathrm{rad}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$, but also the corresponding photon number densities, $N^{\mathrm{R}}_{\mathrm{rad}}$ and $N^{\mathrm{N}}_{\mathrm{rad}}$. A similar approach has been adopted in several studies in numerical relativity, where both energy and number densities are evolved to reconstruct the neutrino spectrum (F. Foucart et al., 2016; D. Radice et al., 2022; C. Musolino & L. Rezzolla, 2024). However, this method has not yet been applied to RT. Analogous to the equations for the radiation energy density, we assume that the photon number density evolves according to equations given in Appendix C.
@@ 4.3 Spectral reconstruction and opacity estimation from mean photon energy
$\langle h\nu\rangle=\frac{f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}}{f_{c}N^{\mathrm{R}}_{\mathrm{rad}}+N^{\mathrm{N}}_{\mathrm{rad}}}.$ (61)
Note that we use the physical radiation energy $f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}$, rather than the conserved quantity $E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}$. We then introduce an effective radiation temperature $T_{\mathrm{rad}}$, which is defined consistently with the mean photon energy as
@@
$\chi_{0F}=\frac{3(1-\xi_{\chi})}{2}\chi_{\mathrm{0R}}+\frac{3\xi_{\chi}-1}{2}\chi_{0E},$ (64)
$\xi_{\chi}=\frac{3+4[|\bm{F}_{S}|/\{c(f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}})\}]^{2}}{5+2\sqrt{4-3[|\bm{F}_{S}|/\{c(f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}})\}]^{2}}},$ (65)
@@
5 Computational Procedure
We implement the RHD scheme described in Section 4 into the self-gravitating magnetohydrodynamics code with adaptive mesh refinement, SFUMATO (T. Matsumoto, 2007). Several updated versions of SFUMATO have been developed to incorporate RT, including SFUMATO-RT (K. Sugimura et al., 2020), which is based on ray tracing, and SFUMATO-M1 (H. Fukushima & H. Yajima, 2021), which employs the M1 closure method combined with the RSLA. These extensions have been applied to a variety of astrophysical simulations, and the implementation presented here further broadens the code’s applicability.
We solve Equations (1), (53)–(59), and (C1)–(LABEL:eq:Gprime), together with the Poisson equation, using an operator-splitting method at each time step. We decompose the equations into the following five parts and solve them in order. In what follows, we describe the solution procedure for each step in some detail. The treatment of the photon number density is identical to the radiation energy density and omitted here for brevity.
5.1 Step 1: hydrodynamics step
@@
Here, we evaluate the intercell numerical fluxes in Equations (66)–(68) using the HLLC approximate Riemann solver (E. F. Toro et al., 1994). On the other hand, we compute the fluxes in Equation (69) as the product of the interface velocity obtained from the Riemann solver and the $(4/3)(E^{\mathrm{R}}_{\mathrm{rad,T}}+E^{\mathrm{N}}_{\mathrm{rad}})$ in the upwind cell.
In the original SFUMATO code, a monotone upstream-centered scheme for conservation laws (MUSCL) and a predictor-corrector method are employed to achieve second-order accuracy in both space and time (T. Matsumoto, 2007). However, in regions with steep gradients in physical quantities, such as protostellar surfaces, this type of interpolation can yield inaccurate results and lead to breakdown of simulations. To avoid such issues and enhance numerical stability, we adopt a simpler approach with first-order accuracy in both space and time in this study. We estimate the values at cell interfaces by taking them equal to those at cell centers, and solve the time evolution using the Euler method.
5.2 Step 2: gravity step
@@
where $G$ is the gravitational constant, $\phi$ is the gravitational potential, and $\bm{g}=-\nabla\phi$ is the gravitational acceleration. We solve Equation (70) to obtain the gravitational potential $\phi$ using a multigrid method, as described in detail by T. Matsumoto (2007). We then solve Equations (71) and (72) following the method by P. D. Mullen et al. (2021). Unlike conventional self-gravity updates, which treat gravity simply as source terms, their method evaluates the source terms in a way that is mathematically equivalent to flux divergences. This guarantees conservation of total energy and momentum to round-off accuracy. Moreover, a key advantage of their formulation is that, unlike previous energy-conservative approaches, it also guarantees curl-free gravity, $\nabla\times\bm{g}=0$. In simulations of self-gravitating systems, numerical errors that violate either energy–momentum conservation or the curl-free nature of gravity can accumulate and lead to unphysical behavior (Y.-F. Jiang & J. Goodman, 2011; Y.-F. Jiang et al., 2013; P. D. Mullen et al., 2021). Therefore, the energy-conservative self-gravity scheme by P. D. Mullen et al. (2021) is essential for maintaining long-term stability and suppressing spurious numerical artifacts. We examine in Appendix D how the protostellar evolution during the collapse of a spherically symmetric gas depends on the choice of self-gravity scheme.
In the method by P. D. Mullen et al. (2021), the right-hand side of Equation (71) is evaluated with the mass flux obtained from the Riemann solver. This treatment ensures that the energy source term is consistent with mass conservation. Following this prescription, we compute the right-hand side of Equation (71) using the Riemann mass flux obtained in Step 1 (hydrodynamics step). Although their implementation employs a predictor–corrector time integrator, we instead use a first-order time integration with the Euler method to enhance numerical stability.
5.3 Step 3: emission and absorption step
@@
Here, we distribute $E^{\mathrm{con}}_{\mathrm{rad}}$ into $E^{\mathrm{R}}_{\mathrm{rad}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$ following Equation (49). As described in Section 4.2, to accurately capture the escape of radiation from an optically thick cell into an optically thin one, we must take into account information from neighboring cells and estimate $\eta^{\mathrm{N}}$ using the minimum cell optical depth among the local and neighboring cells. Accordingly, we compute $\eta^{\mathrm{N}}$ and distribute $E^{\mathrm{con}}_{\mathrm{rad}}$ as
$\tau^{i,j,k}_{\mathrm{cell,eff}}=\min_{(i^{\prime},j^{\prime},k^{\prime})\in\mathcal{N}(i,j,k)}\tau^{i^{\prime},j^{\prime},k^{\prime}}_{\mathrm{cell}},$ (76)
@@
where the superscripts $(i,j,k)$ denote the cell indices in the 3D computational grid, and $\mathcal{N}(i,j,k)$ represents the set consisting of the cell $(i,j,k)$ and its eight neighboring cells.
@@
5.4 Step 4: radiative transfer step

### v2

@@ IV.1 Hybrid radiative transfer: RSLA and Non-RSLA components
As discussed in Section III.1, applying the RSLA violates physical energy conservation, making it difficult to accurately follow the time evolution of optically thick systems, especially when the radiation energy dominates over the gas internal energy. However, in very optically thick regions, radiation propagates slowly via diffusion, and therefore it is possible to evolve the radiation without applying the RSLA while maintaining a reasonably large time step. Indeed, considering the diffusion approximation given in Equation (13), von Neumann stability analysis requires the time step to satisfy
@@
Motivated by this, we introduce an additional radiation energy component, denoted as $E^{\mathrm{N}}_{\mathrm{rad}}$, which is evolved without the RSLA and dominates in regions where $\chi_{\mathrm{0R}}\geq\chi_{\mathrm{crit}}$. Before presenting the detailed formulation, we first summarize the three radiation components used throughout this paper. First, we consider two components to which the RSLA is applied:
(1) RSLA–streaming component, $E^{\mathrm{R}}_{\mathrm{rad,S}}$. — This component represents the portion of the radiation field subject to the RSLA that is transported using the standard M1 closure.
(2) RSLA–trapped component, $E^{\mathrm{R}}_{\mathrm{rad,T}}$. — This component corresponds to the remaining part of the RSLA radiation field that is not transported by the M1 scheme.
The total RSLA radiation energy is given by $E^{\mathrm{R}}_{\mathrm{rad}}=E^{\mathrm{R}}_{\mathrm{rad,S}}+E^{\mathrm{R}}_{\mathrm{rad,T}}$. As noted in Section II.3, the conserved quantity of the RSLA component in the gas–radiation interaction is the radiation energy density divided by the reduction factor $f_{c}$, as defined in Equation (35). The $E^{\mathrm{R}}_{\mathrm{rad}}$ corresponds to this conserved quantity. Next, we consider the third component, to which the RSLA is not applied:
(3) non-RSLA component, $E^{\mathrm{N}}_{\mathrm{rad}}$. — This component is transported diffusively with the true speed of light.
Because the non-RSLA component evolves at the full speed of light, its radiation energy $E^{\mathrm{N}}_{\mathrm{rad}}$ is the conserved quantity. Therefore, in the region where $\chi_{\mathrm{0R}}\geq\chi_{\mathrm{crit}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$ dominates, we can accurately follow the evolution of optically thick systems while conserving energy.
@@
during the calculation. The timing of this redistribution can be chosen appropriately depending on the specific implementation of the code (see Section V). With this formulation, if an optically thick system contains regions where $\tau_{\mathrm{cell}}<2/(3f_{c})$, local energy conservation breaks down in such regions, since the RSLA component dominates. We discuss this impact in Section VIII.1.
@@
where $\bm{G}^{\mathrm{R}}_{\mathrm{S}}$ is the radiation force associated with the RSLA streaming component, $\bm{G}^{\mathrm{R}}_{\mathrm{T}}$ the radiation-pressure gradient arising from the RSLA trapped component, and $\bm{G}^{\mathrm{N}}$ the radiation-pressure gradient from the non-RSLA diffusive component. The total spatial component of radiation four-force density is given by the sum of these contributions, $\bm{G}_{\mathrm{tot}}=\bm{G}^{\mathrm{R}}_{\mathrm{S}}+\bm{G}^{\mathrm{R}}_{\mathrm{T}}+\bm{G}^{\mathrm{N}}$. Note that we do not evolve the three radiation components, $E^{\mathrm{R}}_{\mathrm{rad,S}}$, $E^{\mathrm{R}}_{\mathrm{rad,T}}$, and $E^{\mathrm{N}}_{\mathrm{rad}}$, individually, but rather their sum, $E_{\rm rad}^{\rm con}$ with Equation (55). From Equations (53) and (55), we can see that the $E_{\mathrm{gas}}+E^{\mathrm{con}}_{\mathrm{rad}}$ is conserved during the gas-radiation interaction. We solve Equations (53)–(59) in combination with Equations (1), (36), (44), (46), and (49).
IV.2 Radiation distribution based on neighboring cell information
We distribute the conserved quantity $E^{\mathrm{con}}_{\mathrm{rad}}$ among $E^{\mathrm{R}}_{\mathrm{rad,S}}$, $E^{\mathrm{R}}_{\mathrm{rad,T}}$, and $E^{\mathrm{N}}_{\mathrm{rad}}$ based on the optical depth per cell $\tau_{\mathrm{cell}}$, as described in Equations (16) and (49). However, as mentioned in Section III.2, this approach fails to capture the leakage of radiation from optically thick cells into neighboring optically thin cells, because the radiation energy in an optically thick cell is stored in either $E^{\mathrm{R}}_{\mathrm{rad,T}}$ or $E^{\mathrm{N}}_{\mathrm{rad}}$, neither of which is allowed to freely leave the cell. As a result, we cannot capture the outgoing radiation flux from optically thick cell surfaces. This issue is particularly relevant to the narrow layers near the stellar surface, where the gas density and opacity rise steeply.
@@ IV.2 Radiation distribution based on neighboring cell information
which approximately reproduces the physically expected surface flux. We describe in Section V the details of when and how the optical depths of neighboring cells are incorporated into the calculation, together with the code implementation.
IV.3 Spectral reconstruction and opacity estimation from mean photon energy
As mentioned in Section III.3, to accurately capture the interaction between the gas and radiation, we need to know the radiation spectrum. However, the radiation energy density alone does not uniquely determine the spectral shape. To address this issue, we evolve not only the radiation energy densities, $E^{\mathrm{R}}_{\mathrm{rad}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$, but also the corresponding photon number densities, $N^{\mathrm{R}}_{\mathrm{rad}}$ and $N^{\mathrm{N}}_{\mathrm{rad}}$. A similar approach has been adopted in several studies in numerical relativity, where both energy and number densities are evolved to reconstruct the neutrino spectrum (Foucart et al., 2016; Radice et al., 2022; Musolino and Rezzolla, 2024). However, this method has not yet been applied to RT. Analogous to the equations for the radiation energy density, we assume that the photon number density evolves according to equations given in Appendix C.
@@ IV.3 Spectral reconstruction and opacity estimation from mean photon energy
$\langle h\nu\rangle=\frac{f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}}{f_{c}N^{\mathrm{R}}_{\mathrm{rad}}+N^{\mathrm{N}}_{\mathrm{rad}}},$ (61)
where $h$ is the Planck constant. Note that we use the physical radiation energy $f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}$, rather than the conserved quantity $E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}$. We then introduce an effective radiation temperature $T_{\mathrm{rad}}$, which is defined consistently with the mean photon energy as
@@
$\begin{array}[]{l}{\displaystyle\chi_{0F}=\frac{3(1-\xi_{\chi})}{2}\chi_{\mathrm{0R}}+\frac{3\xi_{\chi}-1}{2}\chi_{0E},}\\[10.0pt] {\displaystyle\xi_{\chi}=\frac{3+4[|\bm{F}_{S}|/\{c(f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}})\}]^{2}}{5+2\sqrt{4-3[|\bm{F}_{S}|/\{c(f_{c}E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}})\}]^{2}}},}\end{array}$ (66)
@@
For clarity, we summarize the opacities used in the RHD equations. Energy exchange between gas and radiation (Equation 57) uses the Planck mean opacity $\kappa_{\mathrm{0P}}$ and the energy mean opacity $\kappa_{0E}$ given by Equations (8) and (9). Momentum exchange between the gas and the RSLA streaming component (Equation 58) uses the flux-mean opacity $\chi_{0F}$, whose behavior transitions smoothly according to the optical depth following Equation (66). In the optically thin limit, $\chi_{0F}$ reduces to the energy mean opacity $\chi_{0E}$ (Equation 63), whereas in the optically thick diffusion limit it approaches the Rosseland mean opacity $\chi_{0\mathrm{R}}$ (Equation 12), consistent with the diffusion approximation. The quantity $\hat{\chi}$ appearing on the left-hand side of Equation (55) governs the transport of the non-RSLA component in optically thick regions, and is therefore evaluated using the Rosseland mean opacity, as defined in Equation (51). The corresponding opacities for photon-number evolution are computed in an analogous manner.
Finally, we note that the present spectral reconstruction implicitly assumes that the radiation field in each cell can be represented by a single-temperature, Planck-like spectrum. The method accurately reproduces spectral shapes, including blackbody emission from the protostar and thermal emission from dust. By self-consistently tracking the mean photon energy in space and time, the scheme enables radiative feedback processes, such as heating and radiation pressure, to be evaluated with good accuracy.
V Computational Procedure
We implement the RHD scheme described in Section IV into the self-gravitating magnetohydrodynamics code with adaptive mesh refinement, SFUMATO (Matsumoto, 2007). Several updated versions of SFUMATO have been developed to incorporate RT, including SFUMATO-RT (Sugimura et al., 2020), which is based on ray tracing, and SFUMATO-M1 (Fukushima and Yajima, 2021), which employs the M1 closure method combined with the RSLA. These extensions have been applied to a variety of astrophysical simulations, and the implementation presented here further broadens the code’s applicability.
We solve Equations (1), (53)–(59), and (C1)–(C4), together with the Poisson equation, using an operator-splitting method at each time step. We decompose the equations into the following five parts and solve them in order. In what follows, we describe the solution procedure for each step in some detail. Because the density and temperature of each cell evolve during every substep, the optical depth of a cell also varies accordingly. Thus, the redistribution of the radiation components based on Equations (16) and (49) is also performed at every substep in which the effects of radiation are involved. The detailed procedure for this redistribution is described below. The treatment of the photon number density is identical to the radiation energy density and omitted here for brevity.
V.1 Step 1: hydrodynamics step
@@
Here, we evaluate the intercell numerical fluxes in Equations (67)–(69) using the HLLC approximate Riemann solver (Toro et al., 1994). On the other hand, we compute the fluxes in Equation (70) as the product of the interface velocity obtained from the Riemann solver and the $(4/3)(E^{\mathrm{R}}_{\mathrm{rad,T}}+E^{\mathrm{N}}_{\mathrm{rad}})$ in the upwind cell. Here, we estimate $E^{\mathrm{R}}_{\mathrm{rad,T}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$ using the local $\tau_{\mathrm{cell}}$ with Equation (16) and (49).
In the original SFUMATO code, a monotone upstream-centered scheme for conservation laws (MUSCL) and a predictor-corrector method are employed to achieve second-order accuracy in both space and time (Matsumoto, 2007). However, in regions with steep gradients in physical quantities, such as protostellar surfaces, this type of interpolation can yield inaccurate results and lead to breakdown of simulations. To avoid such issues and enhance numerical stability, we adopt a simpler approach with first-order accuracy in both space and time in this study. We estimate the values at cell interfaces by taking them equal to those at cell centers, and solve the time evolution using the Euler method.
V.2 Step 2: gravity step
@@
where $G$ is the gravitational constant, $\phi$ is the gravitational potential, and $\bm{g}=-\nabla\phi$ is the gravitational acceleration. We solve Equation (71) to obtain the gravitational potential $\phi$ using a multigrid method, as described in detail by Matsumoto (2007). We then solve Equations (72) and (73) following the method by Mullen et al. (2021). Unlike conventional self-gravity updates, which treat gravity simply as source terms, their method evaluates the source terms in a way that is mathematically equivalent to flux divergences. This guarantees conservation of total energy and momentum to round-off accuracy. Moreover, a key advantage of their formulation is that, unlike previous energy-conservative approaches, it also guarantees curl-free gravity, $\nabla\times\bm{g}=0$. In simulations of self-gravitating systems, numerical errors that violate either energy–momentum conservation or the curl-free nature of gravity can accumulate and lead to unphysical behavior (Jiang and Goodman, 2011; Jiang et al., 2013; Mullen et al., 2021). Therefore, the energy-conservative self-gravity scheme by Mullen et al. (2021) is essential for maintaining long-term stability and suppressing spurious numerical artifacts. We examine in Appendix D how the protostellar evolution during the collapse of a spherically symmetric gas depends on the choice of self-gravity scheme.
In the method by Mullen et al. (2021), the right-hand side of Equation (72) is evaluated with the mass flux obtained from the Riemann solver. This treatment ensures that the energy source term is consistent with mass conservation. Following this prescription, we compute the right-hand side of Equation (72) using the Riemann mass flux obtained in Step 1 (hydrodynamics step). Although their implementation employs a predictor–corrector time integrator, we instead use a first-order time integration with the Euler method to enhance numerical stability.
V.3 Step 3: emission and absorption step
@@
Here, we distribute $E^{\mathrm{con}}_{\mathrm{rad}}$ into $E^{\mathrm{R}}_{\mathrm{rad}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$ following Equation (49). As described in Section IV.2, to accurately capture the escape of radiation from an optically thick cell into an optically thin one, we must take into account information from neighboring cells and estimate $\eta^{\mathrm{N}}$ using the minimum cell optical depth among the local and neighboring cells. Accordingly, we compute $\eta^{\mathrm{N}}$ and distribute $E^{\mathrm{con}}_{\mathrm{rad}}$ as
$\tau^{i,j,k}_{\mathrm{cell,eff}}=\min_{(p,q,r)\in\{(i,j,k)\}\cup\mathcal{N}(i,j,k)}\tau^{p,q,r}_{\mathrm{cell}},$ (77)
@@
where the superscripts $(i,j,k)$ and $(p,q,r)$ denote cell indices in the three-dimensional computational grid, and $\mathcal{N}(i,j,k)$ denotes the set of the six face-adjacent neighboring cells of cell $(i,j,k)$. The use of the six face-adjacent neighboring cells in evaluating $\tau_{\mathrm{cell,eff}}$ is motivated by the radiative transfer step (Section V.4), in which radiative transport is solved exclusively between a cell and its six neighboring cells. If any of these neighboring cells is optically thin, radiation should preferentially escape in that direction. By defining $\tau_{\mathrm{cell,eff}}$ as above, we ensure that sufficient radiation is assigned to the RSLA streaming component when an optically thin escape channel exists.
@@
V.4 Step 4: radiative transfer step